Albuterold, a bronchodilator of the beta-2 adrenergic drug type, can be administered by nebulizer or inhaler. A nebulizer trwatnebf for an adult patient consists of 2.5 mg of albuterol in the form of albuterol sulfate soltion. This is administered for 5 to 15 minutes, three to four times a day. If the concentration of the solution is 1.25 mg/3 ml, how many milliliters are required for four administrations?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required for four administration.

        Formula

             Volume(ml) = Desired dose/Dose in hand * Quantity.

    Here

         Desired dose (D) = 2.5 mg (four times a day).

         Dose in hand (H) = 1.25 mg.

              Quantity (Q) = 3 ml

As per above formula

         Volume of drug required = D/H * Q

                                                = 2.5/1.25 * 3

                                                = 2 * 3

                                                = 6 ml (four times a day).

    Volume(ml) of drug required for four administration = 6 * 4

                                                                                       = 24 ml.